I am trying to work out this problem. Please check my work.What rule should apply in this one? Thanks to all of you.
Your state government decides to raise money by running a lottery where each tickets cost $1.00 and you must choose four different whole numbers between 1-20. When 4 numbers between 1-20 are randomly selected, what is the probability of winning if you have one ticket and the winning numbers can be in any order?
4/20 * 4/4 = 16/80 = 1/5 or 0.2

Answer by stanbon(48568)

Since the winning numbers can be in any order there are 4! winning patterns.
Since there are 20 numbers to choose from there are 20C4 patterns.
Prob of having a winning pattern = 4!/(20C4)=24/4845=8/1615=0.00495....
Cheers,
Stan H.